# Efficient Bayesian Additive Regression Models For Microbiome Studies

### v1

@@ 6 Discussion
Motivated by applied statistical challenges in the analysis of microbiota data, we have introduced a class of Bayesian Multinomial Logistic-Normal additive linear and non-linear regression models (MultiAddGPs). Using recent theory on Marginally Latent Matrix-t Processes (MLTPs), we have developed efficient inference for this class of models. Beyond MultiAddGPs we also introduced a general framework for hyperparameter selection in MLTPs via Maximum Marginal Likelihood (MML) and penalized MML using the Laplace Approximation within the Collapse-Uncollapse (CU) sampler. These models and methods have been made widely available as part of the fido R package on CRAN. Through both simulated and real data analyses, we have shown that this class of models provides a practical, flexible, and rigorous approach to modeling microbiota data. Still, key challenges and future directions remain.
@@
This article develops an efficient approach to hyperparameter selection via approximate MML and penalized MML estimation. An advantage of our approach is its flexibility: researchers can optimize hyperparameters within a kernel function or even within the log-ratio transformation. In practice, we have found Bayesian Optimization to be a practical approach to maximizing the approximate marginal likelihood with small to moderate numbers of hyperparameters (e.g., ¡10). Further research and improvement are likely needed when selecting larger numbers of hyperparameters simultaneously or when the optimization landscape is particularly challenging.
@@
Besides the challenge of calculating MML and penalized MML estimates, alternative methods of selecting hyperparameters may be needed. In particular, some researchers may require fully Bayesian solutions that can quantify uncertainty in model hyperparameters. This is the notable advantage of Cheng et al. (2019) approach. While their model ignores counting variation, and assumes the data is transformed Gaussian, their method provides fully Bayesian inference of hyperparameters via the slice sampler of Murray and Adams (2010). We expect that the same slice-sampling approach could be used for MultiAddGP, replacing the Metropolis Hastings Steps used in that method with the CU sampler with marginal Laplace approximation.
@@
Finally, readers should be aware of the limits of the Laplace Approximation in the CU sampler when inferring MLN-MLTPs. In Silverman et al. (2022), we proved that the error rate of this approximation is of order $O_{p}([D-1]\sum_{n=1}^{N}[\sum_{d=1}^{D}Y_{dn}]^{-1})$. This result implies that the approximation works well when the average number of counts per sample (e.g., the sequencing depth) is high. This is the case with modern sequencing technologies. In contrast, we expect the approximation will be poor when there are few counts per sample, for example, when observations are categorical ($\sum_{d=1}^{D}Y_{dn}=1$). In the latter case, other approaches are needed.
{acks}
[Acknowledgments] The authors thank Rachel Silverman and Hangzhi Guo for their helpful comments. {funding} TC, MPN, and JDS were supported in part by NIH Grant 1 R01 GM148972-01.
[…]

### v2

@@ 6 Discussion
Motivated by applied statistical challenges in the analysis of microbiota data, we have introduced a class of Bayesian Multinomial Logistic-Normal additive linear and non-linear regression models (MultiAddGPs). Using recent theory on Marginally Latent Matrix-t Processes (MLTPs), we have developed efficient inference and hyperparameter estimation algorithms for this class of models. Additionally, we have incorporated non-stationary kernel functions designed to model treatment interventions and disease effects, allowing capture of complex microbiome dynamics. Beyond MultiAddGPs, our hyperparameter estimation techniques can be applied in the much larger class of MLTPs that can be sampled using the CU Sampler. Our methods have been made widely available as part of the fido R package, which is published on CRAN.
While our methods are inspired by challenges in microbiome data, MultiAddGPs could be applied broadly to other settings where count compositional data is collected, including gene expression studies (Fernandes et al. 2014), macro-ecology (Billheimer et al. 2001), and even political polling (Topîrceanu & Precup 2020). Still, key challenges and future directions remain.
@@
This article develops an efficient approach to hyperparameter selection via approximate MML and penalized MML estimation. An advantage of our approach is its flexibility: researchers can optimize hyperparameters within a kernel function or even within the log-ratio transformation. In practice, we have found Bayesian Optimization to be a practical approach to maximizing the approximate marginal likelihood with small to moderate numbers of hyperparameters (e.g., <10). Further research and improvement are likely needed when selecting larger numbers of hyperparameters simultaneously or when the optimization landscape is particularly challenging.
@@
Besides the challenge of calculating MML and penalized MML estimates, alternative methods of selecting hyperparameters may be needed. In particular, some researchers may require fully Bayesian solutions that can quantify uncertainty in model hyperparameters. This is the notable advantage of Cheng et al. (2019) approach. While their model ignores counting variation, and assumes the data is transformed Gaussian, their method provides fully Bayesian inference of hyperparameters via the slice sampler of Murray & Adams (2010). We expect that the same slice-sampling approach could be used for MultiAddGPs, replacing the Metropolis-Hastings Steps used in that method with the CU sampler with marginal Laplace approximation.
@@
Finally, readers should be aware of the limits of the Laplace Approximation in the CU sampler when inferring Multinomial Logistic-Normal MLTPs. In Silverman et al. (2022), we proved that the error rate of this Laplace approximation is of order $O_{p}([D-1]\sum_{n=1}^{N}[\sum_{d=1}^{D}Y_{dn}]^{-1})$. This result implies that the approximation works well when the average number of counts per sample (e.g., the sequencing depth) is high. This is the case with modern sequencing technologies. In contrast, we expect the approximation will be poor when there are few counts per sample, for example, when observations are categorical ($\sum_{d=1}^{D}Y_{dn}=1$). In this scenario, alternative methods, such as the Debiased Multinomial Dirichlet Bootstrap from  Saxena et al. (2024), may be necessary.
7 Disclosure statement
The authors have the following conflicts of interest to declare.
8 Data Availability Statement
All data and code needed to reproduce the analyses in this article are available at https://github.com/Silverman-Lab/MultiAddGPs.
SUPPLEMENTAL MATERIALS
Supplement
includes the following parts: a review of the Collapse-Uncollapsed (CU) sampler and Laplace Approximation in the Collapsed step, a derivation of the posterior distribution for the matrix conjugate linear model, a theoretical result on MLTP for MultiAddGP models, and additional simulation and application details.